# SLM-Mod: Small Language Models Surpass LLMs at Content Moderation

### v1

@@ Appendix B LoRA Hyperparameters
In order to fine-tune our community-specific LLMs, we perform Low-Rank Adaptation (LoRA) Hu et al. (2021) for 1 epoch on 8000 balanced samples from publicly available content moderation datasets. We use rank $r=16$ LoRA with an $\alpha=32$ and no dropout. We use $5$ warmup steps, a linear learning rate schedule with $lr=2e-4$, AdamW Loshchilov (2017) optimizer with a weight decay of $0.01$.
@@ Appendix D Detailed Cross-Domain Results
Figure 5: Cross-domain Moderation Performance for Llama-3.1-8b
Figure 6: Cross-domain Moderation Performance for Gemma-2-9b
Figure 7: Cross-domain Moderation Performance for Mistral-NeMo-Instruct
Appendix E Detailed Cross-domain Correlation Test Results
In this section, we provide a detailed set of results from the Pearson Correlation Coefficient t-test for non-correlation that we conducted to identify any patterns in cross-domain model performance. Table 6 shows test results for relative subreddit size, Table 7 shows the test results for embeddings of subreddit descriptions, and Table 8 shows test results for embeddings of subreddit rules.
@@ Appendix E Detailed Cross-domain Correlation Test Results
llama-3.1-8b gemma-2-9b mistral-nemo-instruct
AskHistorians -0.068 -0.331 -0.180
AskReddit 0.010 -0.033 0.099
AskWomen 0.256 0.153 0.275
Futurology -0.340 -0.421 -0.388
Games -0.218 -0.280 -0.225
IAmA 0.033 -0.106 -0.036
anime 0.102 -0.138 0.045
askscience -0.026 -0.231 -0.190
aww 0.073 0.063 0.033
changemyview -0.151 -0.108 -0.314
history -0.319 -0.309 -0.143
movies -0.029 -0.003 -0.080
nba 0.296 0.259 0.295
politics 0.062 -0.033 -0.101
science -0.309 -0.287 -0.291
@@
llama-3.1-8b gemma-2-9b mistral-nemo-instruct
AskHistorians 0.459 ($\ast$) 0.370 0.264
AskReddit 0.009 -0.063 -0.037
AskWomen -0.074 0.030 -0.078
Futurology 0.099 0.152 0.116
Games 0.196 0.224 0.195
IAmA -0.004 0.045 -0.020
anime 0.160 0.172 0.135
askscience 0.314 0.536 ($\ast$) 0.329
aww -0.070 0.138 0.042
changemyview 0.166 0.114 0.131
history 0.284 0.018 -0.046
movies -0.168 -0.086 -0.139
nba -0.073 -0.264 -0.179
politics -0.155 -0.202 -0.371
science 0.377 0.273 0.367
@@
llama-3.1-8b gemma-2-9b mistral-nemo-instruct
AskHistorians 0.009 0.008 -0.103
AskReddit 0.173 -0.007 0.308
AskWomen 0.307 0.244 0.298
Futurology -0.116 -0.058 -0.057
Games 0.004 -0.084 -0.102
IAmA 0.025 -0.021 0.065
anime 0.266 0.459 ($\ast$) 0.357
askscience -0.223 0.020 -0.361
aww 0.257 0.201 0.229
changemyview 0.134 -0.029 -0.018
history 0.077 -0.327 -0.337
movies 0.344 0.262 0.322
nba 0.462 ($\ast$) 0.439 0.388
politics 0.416 0.454 0.330
science -0.054 -0.007 -0.010

### v2

@@ Appendix B LoRA Hyperparameters
In order to fine-tune our community-specific LLMs, we perform Low-Rank Adaptation (LoRA) Hu et al. (2021) for 1 epoch on 8,000 balanced samples from publicly available content moderation datasets. We use rank $r=16$ LoRA with an $\alpha=32$ and no dropout. We use $5$ warmup steps, a linear learning rate schedule with $lr=2e-4$, AdamW Loshchilov (2017) optimizer with a weight decay of $0.01$.
@@ Appendix D Detailed Cross-Domain Results
Figure 6: Cross-domain Moderation Performance for Llama-3.1-8b.
Figure 7: Cross-domain Moderation Performance for Gemma-2-9b.
Figure 8: Cross-domain Moderation Performance for Mistral-NeMo-Instruct.
Appendix E Performance of Base SLMs on Content Moderation Task Without Fine-tuning
In this section, we report the performance of base SLMs without fine-tuning on the in-domain content moderation tasks described in the main text.
From Table 3 we observe that SLMs are ill-suited for content moderation prior to fine-tuning as compared to LLMs, which is expected given the difference in model sizes, capabilities, and lack of instruction-tuning. As a result, we decided to pursue fine-tuned SLMs for this work since they are more interesting to study.
We observe that being an instruction-tuned model, Mistral-NeMo-Instruct performs the best and provides acceptable performance even prior to fine-tuning, but the performance of Llama-3.1-8b and Gemma-2-9b are near random chance.
Appendix F Detailed Cross-domain Correlation Test Results
In this section, we provide a detailed set of results from the Pearson Correlation Coefficient t-test for non-correlation that we conducted to identify any patterns in cross-domain model performance. Table 8 shows test results for relative subreddit size, Table 9 shows the test results for embeddings of subreddit descriptions, and Table 10 shows test results for embeddings of subreddit rules. Overall, we find only four instances of statistically significant correlation across all subreddits and models.
@@ Appendix F Detailed Cross-domain Correlation Test Results
Subreddit llama-3.1-8b gemma-2-9b mistral-nemo-instruct
r/AskHistorians -0.068 -0.331 -0.180
r/AskReddit 0.010 -0.033 0.099
r/AskWomen 0.256 0.153 0.275
r/Futurology -0.340 -0.421 -0.388
r/Games -0.218 -0.280 -0.225
r/IAmA 0.033 -0.106 -0.036
r/anime 0.102 -0.138 0.045
r/askscience -0.026 -0.231 -0.190
r/aww 0.073 0.063 0.033
r/changemyview -0.151 -0.108 -0.314
r/history -0.319 -0.309 -0.143
r/movies -0.029 -0.003 -0.080
r/nba 0.296 0.259 0.295
r/politics 0.062 -0.033 -0.101
r/science -0.309 -0.287 -0.291
@@
Subreddit llama-3.1-8b gemma-2-9b mistral-nemo-instruct
r/AskHistorians 0.459 ($\ast$) 0.370 0.264
r/AskReddit 0.009 -0.063 -0.037
r/AskWomen -0.074 0.030 -0.078
r/Futurology 0.099 0.152 0.116
r/Games 0.196 0.224 0.195
r/IAmA -0.004 0.045 -0.020
r/anime 0.160 0.172 0.135
r/askscience 0.314 0.536 ($\ast$) 0.329
r/aww -0.070 0.138 0.042
r/changemyview 0.166 0.114 0.131
r/history 0.284 0.018 -0.046
r/movies -0.168 -0.086 -0.139
r/nba -0.073 -0.264 -0.179
r/politics -0.155 -0.202 -0.371
r/science 0.377 0.273 0.367
@@
Subreddit llama-3.1-8b gemma-2-9b mistral-nemo-instruct
r/AskHistorians 0.009 0.008 -0.103
r/AskReddit 0.173 -0.007 0.308
r/AskWomen 0.307 0.244 0.298
r/Futurology -0.116 -0.058 -0.057
r/Games 0.004 -0.084 -0.102
r/IAmA 0.025 -0.021 0.065
r/anime 0.266 0.459 ($\ast$) 0.357
r/askscience -0.223 0.020 -0.361
r/aww 0.257 0.201 0.229
r/changemyview 0.134 -0.029 -0.018
r/history 0.077 -0.327 -0.337
r/movies 0.344 0.262 0.322
r/nba 0.462 ($\ast$) 0.439 0.388
r/politics 0.416 0.454 0.330
r/science -0.054 -0.007 -0.010